How many ounces of pure oil should be mixed with an oil-gas solution that is 3% oil to get 64 ounces of an oil-gas mixture that is 6% oil? Round off to the nearest hundredth of an ounce.
Answer by stanbon(75887)

How many ounces of pure oil should be mixed with an oil-gas solution that is 3% oil to get 64 ounces of an oil-gas mixture that is 6% oil? Round off to the nearest hundredth of an ounce.
===========
Equation::
oil + oil = oil
1*x + 0.03(64-x) = 0.06*64
--------------
100x + 3*64 - 3x = 6*64
---------------------
97x = 9*64
-----
x = 12.54 ounces (amt. of pure oil to add)
---------------
Cheers,
Stan H.
--------------